Is this equation 28 – (4sq rt 2) = 24(sq rt 2) true or false?
Answer by Theo(13342)

by using the calculator, I do not believe they are true.

this makes them false.

the answer I got was:

22.34314575 = 33.9411255

this is clearly false.

your original equation is:

28 - 4*sqrt(2) = 24*sqrt(2)

add 4*sqrt(2) to both sides of the equation to get:

28 = 24*sqrt(2) + 4*sqrt(2)

factor out the sqrt(2) on the right side of the equation to get:

28 = (24 + 4) * sqrt(2) = 28 * sqrt(2)

28 clearly does not equal to 28 * sqrt(2).